OpenStudy (anonymous):

Find all solutions on the interval [0,2π). sin(2r)= - (√3)/2

OpenStudy (anonymous):

I know the rule is sin(U) = c and U = arcsin(c) + k2π or U = π -arcsin(c)+k2π. I'm just kind of trying to figure out how to write it down correctly.

OpenStudy (anonymous):

I have all the answers, but I don't want to break any rules. Obviously I won't be working with anything +k2π, because that would be outside of the interval. I got: r=4π/3 + π or r= 5π/3 +π as my answers. I didn't use the second rule, because they produced answers that weren't equal to -√3/2.

OpenStudy (anonymous):

Would I just write it in set builder notation as {r|r=4π/3 +π and r=5π/6+π}?